1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
Der Werwolf  La lupfantomo  The Banshee (An Approach)
translated by Bertilo Wennergren  translated by Max Knight

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“ Jen lupfantom’ de famili’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

One night, a banshee slunk away

from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der seine Pfoten geduldig kreuzte vor dem Toten:

Do diktis tiu dokta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

�Der Werwolf“ – sprach der gute Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s nennt, den Werwolf, - damit hat’s ein End!“ De “lupfantomo”, simpla rekta, tra “lupfintomo”, plej perfekta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho, koŝmara.

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

Dem Werwolf schmeichelten die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“ Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

The banshee marveled at the cases

and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

La viro ĉagreniĝis tamen:

“Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

…  …  …

2/2

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

http://www.poezio.net/version?poem-id=825&version-id=1652,1653,1660